The Magellanic Clouds Newsletter

An electronic exchange of Magellanic Clouds information

Edited by: You-Hua Chu and Dominik J. Bomans

MCnews@astro.uiuc.edu

http://www.astro.uiuc.edu/mcnews/MCnews.html

Number: 3 March 4, 1996

Table of Contents

News and Views..............1 Paper Abstracts..............2

News and Views

“New Views of the Magellanic Clouds”

A topical session in the 188th AAS meeting at Madison, Wisconsin

We would like to invite you to the topical session “New Views of the Magellanic Clouds” in the 188th American Astronomical Society meeting held at Madison, Wisconsin on June 9-13, 1996. The Magellanic Clouds session, scheduled for June 12 (Wednesday), features seven invited talks on surveys of the Magellanic Clouds and a panel discussion about future research directions and meetings. The invited speakers and their topics include: Smith CCD emission-line line survey of the MCs Staveley-Smith ATCA atomic HI survey of the MCs Kutner SEST CO molecular survey of the MCs Haynes Radio continuum survey of the MCs Snowden ROSAT X-ray survey of the MCs Zaritsky Scanning CCD broad-band survey of the MCs Cook Survey of variable stars (from MACHO project) Contributions to the Magellanic Clouds session should follow the normal AAS abstract submission process. The deadline for abstract submission is April 4, 1996. Please check the WWW page of the AAS at http://www.aas.org/meetings/aas188/prelim/prelim.html for latex templates for abstracts, registration forms, and other information. Please send your abstract to abs-submit@aas.org Contributions can be presented orally or in posters, which needs to be specified in the abstract forms. Tentatively, the poster contributions will be exhibited on June 12, and oral contributions will be scheduled for June 13. If you have any questions, please feel free to contact us. We look forward to seeing you in Madison,

You-Hua Chu, Dominik Bomans, Bart Wakker

Paper Abstracts

The Distance Modulus of the Large Magellanic Cloud: Constraints from RR Lyrae Pulsation Properties

M. Catelan^{1}

1 Instituto Astronômico e Geofísico, Universidade de São Paulo, C.P. 9638, CEP 01065-970, São Paulo, SP, Brasil

It has recently been suggested (van den Bergh 1995, ApJ 451, L65) that the discrepancy between the “long” and “short” distance moduli of the Large Magellanic Cloud (LMC), as inferred from the properties of the Cepheid and RR Lyrae variables, respectively, might be due to the action of “third parameters” between the Galaxy and the LMC, which would make the RR Lyraes in the old LMC globular clusters brighter than their Galactic counterparts by ≃ 0.3 mag. Through analysis of the RR Lyrae pulsation properties, we show that this idea is not supported by the available data. A satisfactory explanation of the problem has yet to be found. Accepted by Astronomy and Astrophysics Letters For preprints, contact marcio@vax.iagusp.usp.br

Be star surveys with CCD photometry. II. NGC 1818 and its neighbouring cluster in the LMC

Eva K. Grebel^{1,2,3}

1 Astronomisches Institut der Universität Würzburg, Am Hubland, D-97074 Würzburg, Germany 2 University of Illinois at Urbana-Champaign, Dept. of Astronomy, 1002 West Green Street, Urbana, IL 61801, USA 3 Sternwarte der Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany

As part of an ongoing photometric survey of young Magellanic Cloud clusters I identified Be stars in NGC 1818 and the neighbouring cluster (21 pc to the southwest) in the Large Magellanic Cloud. The neighbouring cluster does not appear to contain any evolved stars, and its sparsely populated main sequence does not extend to stars as massive as in NGC 1818. Both clusters are younger than the surrounding field population, but the current data do not allow to conclude whether NGC 1818 is a binary cluster or not. The small cluster is more heavily reddened than NGC 1818 indicating the presence of differential reddening or a larger distance for the small cluster.

I find both clusters to contain a large number of Be stars. NGC 1818 contains fewer Be stars than the slightly younger SMC cluster NGC 330, but still comprises about 70 current Be stars down to 18th magnitude. The field only contains very few Be stars as one would expect for a predominantly older population.

Comparing Be star fractions in NGC 1818, NGC 330, and NGC 2004 to young Galactic open clusters, I show that the Be phenomenon is strongly peaked toward early-type B stars (and thus clusters in an age range of a few 10^{7} years). Possible constraints of Be stars and Be star fractions in stellar aggregates of given age and metallicity on Be star theories are discussed. The distribution of Be stars in young clusters suggests that Be stars are in the evolutionary state of main-sequence stars. Submitted to Astronomy & Astrophysics For preprints, contact grebel@astro.uni-wuerzburg.de

Do The Magellanic Cepheids Pose A New Puzzle?

J.R.Buchler^{1}, Z. Kollath^{1}, J.P. Beaulieu^{2} and M.J. Goupil^{3}

1 Physics Department, University of Florida, Gainesville, FL 32611 2 Institut d’Astrophysique, Paris, France 3 DASGAL, Observatoire de Paris, Meudon 92195, France

The observed Magellanic Cloud Cepheid data pose a new challenge to both stellar structure and stellar evolution. The Fourier analysis of the LMC and SMC data indicate that the 2:1 resonance (P_{2}/P_{0} = 0.5) between the fundamental mode of pulsation and the second overtone occurs with a period P_{0} of ≈ 10^{d}, just as in our Galaxy. The implications of this resonance are difficult to reconcile with linear radiative Cepheid models computed for the low metallicity of the MC. They also require a large upward shift in the mass luminosity relation, seemingly larger than is comfortable for evolutionary calculations. The disagreement worsens if we interpret the structure in the first overtone Fourier parameters near 3^{d} as the result of an alleged resonance with the 4^{th} overtone. In contrast to earlier expectations the beat Cepheid data with P_{1}/P_{0} ≈ 0.7 as well as those with P_{2}/P_{1} ≈ 0.8 impose essentially no constraints on the ML relations. Accepted by Astrophysical Journal Letters For preprints, contact buchler@phys.ufl.edu

Results of the ESO-SEST Key Programme: CO in the Magellanic

Clouds V. Further CO observations of the Small Magellanic Cloud

M. Rubio^{1}, J. Lequeux^{2,3}, F. Boulanger^{2,4}, R.S. Booth^{5}, G. Garay^{1}, Th. de Graauw^{6}, F.P. Israel^{7}, L.E.B. Johansson^{5,10}, M.L. Kutner^{8,9} L.–Å. Nyman^{5,10}

1 Departamento de Astronomía , Universidad de Chile, Casilla 36-D, Santiago, Chile 2 Radioastronomie, Ecole Normale Superieure, 24 Rue Lhomond, 75231 Paris CEDEX 05, France 3 DEMIRM, Observatoire de Paris, 61 Av. de l’Observatoire, 75014 Paris, France 4 Institut d’Astrophysique Spatiale, Bat. 120, Université de Paris-XI, 91045 Orsay CEDEX, France 5 Onsala Space Observatory, S-43900 Onsala, Sweden 6 Laboratorium voor Ruimteonderzoek, SRON, Postbus 800, NL-9700 AV Groningen,The Netherlands 7 Sterrewacht, Postbus 9513, NL-2300 Leiden, The Netherlands 8 Department of Physics, Rensselaer Polytechnic Institute, Troy, NY 12180, USA 9 NRAO, 949 N. Cherry Av., Campus Building 65, Tucson , Arizona 85721-0655, USA 10 European Southern Observatory, Casilla 19001, Santiago 19, Chile

We present further results of fully-sampled observations in the J = 1 → 0 and J = 2 → 1 rotational transitions of ^{12}CO and ^{13}CO, obtained with the SEST telescope at 43” or 22” resolution, toward the Small Magellanic Cloud. These observations concern six molecular clouds with a variety of physical conditions. Maps in the^{12}CO(1→0),^{12}CO(2→1), and in some cases of^{13}CO(1→0), lines are presented, as well as profiles of the four lines at the peak positions convolved to the same angular resolution of 43”. The line ratio results are interpreted in Paper IV of this series (Lequeux et al. 1994). Accepted by Astronomy and Astrophysics Supplement Series For preprints, contact mrubio@das.uchile.cl

Planetary Nebulae in the Magellanic Clouds: I) Carbon abundances and Type I nebulae

P. Leisy^{1} and M. Dennefeld^{1}

1 Institut d’Astrophysique (IAP) Paris

We present optical and UV spectroscopy of PNe in the LMC and the SMC. Line intensities in the visible have been measured to a few tenths of percent of Hβ. IUE spectra are used to determine the carbon abundances. This paper discusses the CNO cycle and its implications. We show that the CN or ON cycle are more effective with lower initial metallicities and are always complete for Type I nebulae. The latter cannot easely be distinguished from non-type I PNe on the basis of N or He abundances alone,as continuity exists in all the standard diagrams. From the study of C abundances, we deduce that the 3. dredge-up is occuring in all PNe: its efficiency too is higher with lower initial abundances, offering an easy explanation for the higher number of carbon stars found in metal-deficient galaxies. It is shown that this dredged-up carbon is sometimes transformed into nitrogen by Hot-Bottom Burning, but in a few objects only and not even in all the type I PNe, which therefore do not form a homogeneous class with the standard definition. An alternative one is proposed. Accepted by Astronomy and Astrophysics Supplement Series For preprints, contact dennefel@iap.fr

X-rays from superbubbles in the Large Magellanic Clouds IV: The Blowout Structure of N44

E.A. Magnier^{1}, Y.-H. Chu^{1}, S.D. Points^{1}, U. Hwang^{2}, R.C. Smith^{3}

1 Astronomy Dept., University of Illinois, 1002 W. Green Street, Urbana, IL 61801, USA 2 NASA/Goddard Space Flight Center, Code 666, Greenbelt, MD 20771, USA 3 Dept. of Astronomy, University of Michigan, 834 Dennison, Ann Arbor, MI 48109, USA

We have used optical echelle spectra along with ROSAT and ASCA X-ray spectra to test the hypothesis that the southern portion of the N44 X-ray bright region is the result of a blowout structure. Three pieces of evidence now support this conclusion. First, the filamentary optical morphology corresponding with the location of the X-ray bright South Bar suggests the blowout description (Chu et al. 1993). Second, optical echelle spectra show evidence of high velocity (∼90 km sec^{−1}) gas in the region of the blowout. Third, X-ray spectral fits show a lower temperature for the South Bar than the main superbubble region of Shell 1. Such a blowout can affect the evolution of the superbubble and explain some of the discrepancy discussed by Oey & Massey (1995) between the observed shell diameter and the diameter predicted on the basis of the stellar content and Weaver et al.’s (1977) pressure-driven bubble model. Accepted by Astrophysical Journal For preprints, contact gene@astro.uva.nl http://www.astro.washington.edu/gene/papers/n44.ps.gz